A restaurant chain regularly surveys its customers. On the basis of these surveys, the management of the chain claims that 75% of its customers rate the food as excellent. A consumer testing service wants to examine the claim by asking 460 customers to rate the food.

Seventy per cent rated the food as excellent. Does this support the management’s claim?


1
Expert's answer
2021-03-22T09:58:53-0400

"H_0 : p = 0.75 \\

H_1 : p< 0.75"

σ = 95 %

n = 460

"\bar{x}=0.70 \\

z = \frac{0.70-0.75}{\sqrt{ \frac{0.75 \times 0.25}{460} }} = -2.476 \\

p(-2.4766) = 0.0066"

Since this p-value is less than the level of significance (0.05), the decision is to reject the null hypothesis.